# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v1

@@ 2.6. Global Dimension
With projective dimension as a numeric object, we can directly define global dimension of a ring $R$ as the supremum of the projectve dimensions over all $R$-modulesW.
@@
Basic properties and the relation with HasProjectiveDimensionLE are formalized in the same fileW.(Outside code excerpts we denote globalDimension by $\operatorname{gldim}$)
The supremum over all modules can actually be reduced to the supremum over all finitely generated modules. However, since in mathlib projective dimension is defined via vanishing of Ext and a definition of injective dimension was missing when this formalization started, we formalized only the following weakened form of [4, Lemma 18.2]W.
@@
$(1)\rightarrow(2)$ and $(2)\rightarrow(3)$ are trivial, For $(3)\rightarrow(1)$, the following lemma sufficesW:
Lemma 2.4.
@@
lemma ext_subsingleton_of_quotients [Small.{v} R] (M : ModuleCat.{v} R) (n : ℕ)
(h : I : Ideal R, Subsingleton (Ext (ModuleCat.of R (Shrink.{v} (R I))) M (n + 1))) :
N : ModuleCat.{v} R, Subsingleton (Ext N M (n + 1))
This result is essentially about the $n$-th syzygy of $M$ in an injective resolution. We prove it by induction on $n$:using dimension-shifting on $M$ we reduce to the case $n=0$, which is just an application of the existing Baer criterion. W.
Using Lemma  2.3, we show that the global dimension is the supremum of the projective dimensions of finitely generated modulesW.
@@
Combining this with the localization results from the previous subsection, we formalize the following statementW.
@@
For noetherian ring $R$, $\operatorname{gldim}(R)$ is equal to the supremum of $\operatorname{gldim}(R_{\mathfrak{p}})$ over the prime ideals $\mathfrak{p}$ of $R$.
@@
Since the left-hand side equals the supremum of projective dimensions of finitely generated modules, Lemma  2.1, we know that the left is no greater than the right. Conversely, any $R_{\mathfrak{p}}$-module $N$ is already $R$-module with $N=N_{\mathfrak{p}}$, and Lemma  2.2 yields the reverse inequality, finishing the proof. A version for localizing only at maximal ideals is also providedW.
@@ 3. Definition of Depth and Related Results
Introduction:
In this section we formalize the definition of depth as the minimal index of non-vanishing $\operatorname{Ext}$. Before doing so, and to relate depth to the maximal length of a regular sequence contained in a given ideal, we formalized Rees theorem (see Theorem  3.1). We then establish basic properties of depth, such as the inequalities associated to a short exact sequence and the fact that quotienting by a regular element decreases depth by exactly one. Finally, we present our formalizations of two classical results about depth: the Auslander–Buchsbaum theorem (see 3.2) and Ischebeck’s theorem (see 3.3).
Throughout this section we let $R$ be a Noetherian ring and $M$, NN be $R$-modules.
Theorem 3.1.
(Rees) For $M$ a finitely generated $R$-module, and let $I$ an ideal of $A$ with $IM\neq M$. Fix an integer $n>0$. The following are equivalent.
@@ Introduction:
[4, Theorem 28]
(The equivalence $(2)\longleftrightarrow(4)$ will be a key ingredient in the definition of depth.)
We first prove the following lemma: for finitely generated $M$ and $N$, the $R$-linear maps $N\rightarrow M$ are all zero if and only if there exists a $M$-regular element in the annihilator of MMW.
lemma subsingleton_linearMap_iff [IsNoetherianRing R] [Module.Finite R M] [Module.Finite R N] :
Subsingleton (N →[R] M) r Module.annihilator R N, IsSMulRegular M r
A sketch of the proof : if a prime ideal $\mathfrak{p}\in\mathrm{Ass}(M)$ contains $\mathrm{Ann}(N)$, then the maximal ideal $\mathfrak{m}_{\mathfrak{p}}$ of $R_{\mathfrak{p}}$ is an associated prime of $M_{\mathfrak{p}}$, we can obtain a non-trivial linear map $N_{\mathfrak{p}}\rightarrow N_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}N_{\mathfrak{p}}\rightarrow R_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}$ which yields a nontrivial map $M\rightarrow N$ non-trivial, then applying prime avoidance to $\mathfrak{p}\in\mathrm{Ass}(M)$ finishes the proof.
We now present the proof of the main theorem formalized as followsW.
@@
(M : ModuleCat.{v} R) (Mntr : Nontrivial M) (Mfin : Module.Finite R M) (smul_lt : I · (Top.top : Submodule R M) < Top.top) :
@@
The implications $(1)\rightarrow(2)$ and $(2)\rightarrow(3)$ are straightforward. The implications $(3)\rightarrow(4)$ and $(4)\rightarrow(1)$ follow the arguments in [4, Theorem 28]. The lemma above plays an important role in both directions and is therefore separated as an independent result.
@@ 3.1. Definition of Depth
The Rees theorem above builds a bridge between the homological characterization and the existence of regular sequences. We would like a definition of $\operatorname{depth}$ that is meaningful both in commutative algebra and homological terms. We therefore adopt the following definition (often called the ‘grade’ in some texts)W.
@@
Here $R$ is a commutative ring of Type u that is Small.{v} to ensure ModuleCat.{v} R has enough projectives. These assumptions on $R$ are used throughout in this section. (out side code excerpts, moduleDepth M N will be denoted by $\operatorname{depth}(M,N)$)
@@
We now specialize to the cases of primary interest. The $\operatorname{depth}$ of a module with respect to an ideal is defined as the followingW:
@@
For a local ring, the $\operatorname{depth}$ of a module with respect to the unique maximal ideal isW:
@@
Using Rees theorem, we relate the above homological notion to the maximal length of $M$-regular sequence contained in a given ideal. In the statements below we assume $M$ and $N$ are nontrivial, finitely generated $R$-modules.
If $I$ an ideal of $R$ with $IM<M$ and $\mathrm{Supp}(N)=V(I)$, then $\operatorname{depth}(N,M)=\operatorname{depth}_{I}(M)$W.
lemma moduleDepth_eq_depth_of_supp_eq [IsNoetherianRing R] (I : Ideal R)
(N M : ModuleCat.{v} R) [Module.Finite R M] [Nfin : Module.Finite R N]
(hsupp : Module.support R N = PrimeSpectrum.zeroLocus I) :
moduleDepth N M = I.depth M
Under the same hypotheses, $\operatorname{depth}(N,M)$ equals the supremum of lengths of $M$-regular sequences contained in IIW:
@@
With the homological definition in hand, we obtain the following inequalities for a short exact sequence of nontrivial finitely generated $R$-modules.
Theorem 3.2.
@@
A generalization of [6, 00LX]
These follow readily from the long exact sequence and can be found hereW.
@@
The corresponding dual statements about $\operatorname{depth}(\cdot,M)$ in short exact sequence in short exact sequences are proved similarly using the contravariant long exact sequence, seeW. However for Ideal.depth and IsLocalRing.depth the dual version is rarely used.
We also formalized the behavior of depth under quotient by a regular element.
Theorem $3.$
For a $M$-regular element $x\in\mathrm{Ann}(N)$, we have $\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$
A generalization of [6, 090R]
Indeed, from the short exact sequence $0\rightarrow M\xrightarrow{\cdot x}M\rightarrow M/xM\rightarrow 0$, we obtain long exact sequence
Since $x\in Ann(N)$, the maps $\operatorname{Ext}_{R}^{i}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i}(N,M)$ are zero for all $i$, so $\operatorname{Ext}_{R}^{i}(N,M/xM)$ vanish if and only if both $\operatorname{Ext}_{R}^{i}(N,M)$ and $\operatorname{Ext}_{R}^{i+1}(N,M)$ vanish. From this the theorem followsW.
@@
We also formalized the analogous statement for quotient by a (weakly) regular sequence. This follows by induction on the length of the sequenceW.
lemma moduleDepth_quotient_regular_sequence_add_length_eq_moduleDepth (N M : ModuleCat.{v} R)
(rs : List R) (reg : IsWeaklyRegular M rs) (h : r rs, r Module.annihilator R N) :
moduleDepth N (ModuleCat.of R (M (Ideal.ofList rs) · (Top.top : Submodule R M))) + rs.length = moduleDepth N M
@@ 3.2. The Auslander–Buchsbaum Theorem
In this section we prove the following theorem.
Theorem 3.4.
(Auslander–Buchsbaum) Let $(R,\mathfrak{m},k)$ be a Noetherian local ring and let $M$ be a non-trivial finitely generated $R$-module. If projective dimension of $M$ is finite, then
$\operatorname{projdim}(M)+\operatorname{depth}(M)=\operatorname{depth}(A)$
We first establish a preliminary result. Since we formalized that $\operatorname{Ext}$ commute with finite bi-productW, for a finitely generated free module $M$ over $R$, $\operatorname{Ext}_{R}^{i}(N,M)$ vanish if and only if $\operatorname{Ext}_{R}^{i}(N,R)$ vanish for any $R$-module NNW. As a corollary $\operatorname{depth}(M)=\operatorname{depth}(R)$W.
lemma finte_free_ext_vanish_iff
(M N : ModuleCat.{v} R) [Module.Finite R M] [Module.Free R M] [Nontrivial M] (i : ℕ) :
Subsingleton (Ext.{max u v} N M i) Subsingleton (Ext.{max u v} N (ModuleCat.of R (Shrink.{v} R)) i)
lemma free_depth_eq_ring_depth (M N : ModuleCat.{v} R) [Module.Finite R M] [Module.Free R M] [Nontrivial M] :
moduleDepth N M = moduleDepth N (ModuleCat.of R (Shrink.{v} R))
Back to the main theorem, the case $\operatorname{projdim}(M)=0$ is immediate since then $M$ is projective and hence free.
The case $\operatorname{projdim}(M)=1$ is the core, so we prove it separately.
@@
Pick a $k$ basis of $M/\mathfrak{m}M$ and lift it back to $M$, so we obtain a map $f:R^{n}\rightarrow M$ whose induced map $(R/\mathfrak{m})^{n}\rightarrow M/\mathfrak{m}M$ is isomorphism, hence $\mathrm{ker}(f)\subseteq\mathfrak{m}R^{n}$. Because $\mathrm{ker}(f)$ has projective dimension 0, it is projective thus free. By the following lemma, we know that the inclusion from $\mathrm{ker}(f)$ to $R^{n}$ is contained in $\mathfrak{m}(\mathrm{ker}(f)\rightarrow R^{n})$ (view $R$-linear maps as $R$ module).
theorem mem_smul_top_of_range_le_smul_top [Module.Finite R M] [Module.Free R M] (f : M →[R] N)
(hf : LinearMap.range f I · (Top.top : Submodule R N)) : f I · (Top.top : Submodule R (M →[R] N))
Applying the next lemma with $L=R/\mathfrak{m}$, the map $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})$ induced by the inclusion $\mathrm{ker}(f)\rightarrow R^{n}$ is zero.
lemma ext_hom_zero_of_mem_ideal_smul (L M N : ModuleCat.{v} R) (n : ℕ) (f : M N)
(mem : f (Module.annihilator R L) · (Top.top : Submodule R (M N))) :
(AddCommGrp.ofHom <| ((Ext.mk f)).postcomp L (add_zero n)) = 0
(One easily verifies this lemma by reducing to the case $f=a\cdot g$ with $a\in\mathrm{Ann}(L)$.)
Now consider the long exact sequence induced by $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$.
@@
The first and fourth maps are zero by the discussion above. Moreover, vanishing of $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))$ and of $\operatorname{Ext}_{R}^{i}(k,R^{n})$ are each equivalent to vanishing of $\operatorname{Ext}_{R}^{i}(k,R)$. From the above long exact sequence we deduce that $\operatorname{Ext}_{R}^{i}(k,M)$ vanishes if and only if both $\operatorname{Ext}_{R}^{i}(k,R)$ and $\operatorname{Ext}_{R}^{i+1}(k,R)$ vanish. By the homological characterization of depth this yields $1+\operatorname{depth}(M)=\operatorname{depth}(A)$.
Finally, we prove the theorem by induction on $\operatorname{projdim}(M)$. Assume the statement holds for projective dimension $n\geq 1$, and let $\operatorname{projdim}(M)=n+1$. Take a surjection $f:R^{m}\rightarrow M$, then $\mathrm{ker}(f)$ has projective dimension $n$. By induction hypothesis, $n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$. Since $\operatorname{depth}(R^{m})=\operatorname{depth}(R)\geq n>0$, so there is no non-trivial $k\rightarrow\mathrm{ker}(f)\rightarrow R^{m}$, thus $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$ vanish.
Consider the long exact sequence associated to $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$.
@@
For all $i<\operatorname{depth}(\mathrm{ker}(f))<\operatorname{depth}(R)$, the map $\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))$ is an isomorphism. Together with $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$ vanish, we obtain $\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(M)+1$. Hence $n+1+\operatorname{depth}(M)=n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$, which completes the induction and the proof.

### v2

@@ 2.6. Global Dimension
Recall that the global dimension of a ring is the supremum of projective dimensions over all $R$-modules. With projective dimension as a numeric object, we implement this definition directly.W
@@
Basic properties and the relation with HasProjectiveDimensionLE are formalized in the same fileW.(Outside code excerpts we denote globalDimension by $\operatorname{gldim}$.)
The supremum over all modules can actually be reduced to the supremum over all finitely generated modules. However, since in mathlib projective dimension is defined via vanishing of Ext and a definition of injective dimension was missing when this formalization started, we formalized only the following weakened form of [16, Lemma 18.2].W
@@
Proof.
$(1)\rightarrow(2)$ and $(3)\rightarrow(1)$ are trivial. For $(2)\rightarrow(3)$, the following proposition suffices.W
Proposition 2.12.
@@
Proof.
This result is essentially about the $n$-th syzygy of $M$ in an injective resolution. We prove it by induction on $n$: using dimension-shifting on $M$ we reduce to the case $n=0$, which is just an application of the existing Baer criterionW. ∎
Using Lemma  2.11, we can obtain the following result as a corollary.W
Lemma 2.13 ([8, Theorem 19.1]).
For ring $R$, $\operatorname{gldim}(R)$ is equal to supremum of the projective dimensions of finitely generated modules.
@@
Combining this with the localization results from the previous subsection, we formalize the following statement.W
@@
For a noetherian ring $R$, $\operatorname{gldim}(R)$ is equal to the supremum of $\operatorname{gldim}(R_{\mathfrak{p}})$ over the prime ideals $\mathfrak{p}$ of $R$.
@@
Proof.
Since the left-hand side equals the supremum of projective dimensions of finitely generated modules, by Lemma  2.9, we know that the left is no greater than the right. Conversely, any $R_{\mathfrak{p}}$-module $N$ is already an $R$-module with $N=N_{\mathfrak{p}}$; Lemma  2.10 yields the reverse inequality, completing the proof. ∎
A version for localizing only at maximal ideals is also provided.W
@@ 3. Definition of Depth and Related Results
In this section, we formalize the definition of depth as the minimal index of non-vanishing $\operatorname{Ext}$. First, to relate depth to the maximal length of regular sequences contained in a given ideal, we formalize Rees’ theorem (Theorem 3.1). We then establish basic properties of depth, such as the inequalities associated with short exact sequences and the fact that quotienting by a regular element decreases depth by exactly one. Finally, we present our formalizations of two classical results about depth: the Auslander–Buchsbaum theorem in Subsection  3.2 and Ischebeck’s theorem in Subsection  3.3.
Throughout this section, we let $R$ be a Noetherian ring and $M$, NN be $R$-modules.
Theorem 3.1 (Rees, [16, Theorem 28]).
For $M$ a finitely generated $R$-module, and let $I$ be an ideal of $R$ with $IM\neq M$. Fix an integer $n>0$. The following are equivalent.
@@
(The equivalence $(2)\leftrightarrow(4)$ will be a key ingredient in the definition of depth.)
We now present the proof of the main theorem formalized as follows.W
@@
(M : ModuleCat.{v} R) (Mntr : Nontrivial M) (Mfin : Module.Finite R M)
(smul_lt : I · (Top.top : Submodule R M) < Top.top) :
@@
Proof.
The implications $(1)\rightarrow(2)$ and $(2)\rightarrow(3)$ are straightforward. The implications $(3)\rightarrow(4)$ and $(4)\rightarrow(1)$ follow the arguments in [16, Theorem 28]. ∎
The following proposition plays an important role in both of the two main implications and is therefore separated as an independent result for the convenience of the main proof.W
Proposition 3.2.
For finitely generated $M$ and $N$, the $R$-linear maps $N\rightarrow M$ are all zero if and only if there exists an $M$-regular element in the annihilator of $N$.
Proof.
If a prime ideal $\mathfrak{p}\in\mathrm{Ass}(M)$ contains $\mathrm{Ann}(N)$, then the maximal ideal $\mathfrak{p}R_{\mathfrak{p}}$ of $R_{\mathfrak{p}}$ is an associated prime of $M_{\mathfrak{p}}$. We can obtain a non-trivial linear map $N_{\mathfrak{p}}\rightarrow N_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}N_{\mathfrak{p}}\rightarrow R_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}$ which yields a non-trivial map $M\rightarrow N$. Applying prime avoidance to $\mathfrak{p}\in\mathrm{Ass}(M)$ then completes the proof. ∎
@@ 3.1. Definition of Depth
Rees’ theorem builds a bridge between the homological characterization and the existence of regular sequences. We want a definition of $\operatorname{depth}$ that is meaningful both in commutative algebra and homological terms. We therefore adopt the following definition for the minimal order of non-vanishing $\operatorname{Ext}$ (often called ”grade” in some texts).W
@@
Here $R$ is a commutative ring of Type u that is Small.{v} to ensure ModuleCat.{v} R has enough projectives. These assumptions on $R$ are used throughout in this section. (Out side code excerpts, moduleDepth M N will be denoted by $\operatorname{depth}(M,N)$.)
@@
We now specialize to the cases of primary interest. The $\operatorname{depth}$ of a module with respect to an ideal $I$ is defined by taking the module on the left to be $R/I$, formalized as follows:W
@@
For a local ring, $\operatorname{depth}(M)$ denotes the $\operatorname{depth}$ of $M$ with respect to the unique maximal ideal.W
@@
Using Rees’ theorem, we relate the above homological notion to the maximal length of $M$-regular sequences contained in a given ideal. In the statements below, we assume $M$ and $N$ are non-trivial, finitely generated $R$-modules.
Proposition 3.3.
If $I$ an ideal of $R$ with $IM<M$ and $\mathrm{Supp}(N)=V(I)$, then $\operatorname{depth}(N,M)=\operatorname{depth}_{I}(M)$
Proof.
This is a direct application of the implications $(3)\rightarrow(2)\rightarrow(1)$ of Theorem  3.1. ∎
This result relates the general definition of $\operatorname{depth}(M,N)$ with the usual $\operatorname{depth}_{I}(N)$.W
With the same hypothesis on $I$, we have the following relation with length of regular sequences.W
Proposition 3.4.
$\operatorname{depth}(N,M)$ equals the supremum of lengths of $M$-regular sequences contained in $I$.
@@
Proof.
This is a direct application of the implications $(3)\rightarrow(4)\rightarrow(1)$ of Theorem  3.1. ∎
Returning to the homological definition, we obtain the following inequalities for $\operatorname{depth}$ in a short exact sequence of non-trivial finitely generated $R$-modules.W
Lemma 3.5 (A generalization of [22, 00LX] ).
@@
Proof.
These follow readily from the long exact sequence:
@@
The corresponding ‘dual’ statements about $\operatorname{depth}(\cdot,M)$ in short exact sequences are proved similarly using the contravariant long exact sequence.W However, for Ideal.depth and IsLocalRing.depth, the dual version is rarely used.
We also formalized the behavior of depth under quotient by a regular element.W
Theorem 3.6 (A generalization of [22, 090R]).
For an $M$-regular element $x\in\mathrm{Ann}(N)$, we have $\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$.
@@
Proof.
From the short exact sequence $0\rightarrow M\xrightarrow{\cdot x}M\rightarrow M/xM\rightarrow 0$, we obtain the long exact sequence
Since $x\in Ann(N)$, the maps $\operatorname{Ext}_{R}^{i}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i}(N,M)$ are zero for all $i$, so $\operatorname{Ext}_{R}^{i}(N,M/xM)$ vanish if and only if both $\operatorname{Ext}_{R}^{i}(N,M)$ and $\operatorname{Ext}_{R}^{i+1}(N,M)$ vanish. The theorem follows from this observation. ∎
As a corollary, there is an analogous statement for quotient by a (weakly) regular sequence. W
Corollary 3.7.
For $(r_{1},\cdots,r_{k})$ an $M$-regular sequence contained in $\mathrm{Ann}(N)$, then
$\operatorname{depth}(N,M)+k=\operatorname{depth}(N,M/(r_{1},\cdots,r_{k}))$
Proof.
This result follows easily by induction on the length of the sequence. ∎
@@ 3.2. The Auslander–Buchsbaum Theorem
In this section, we prove the following theorem first introduced in [2].
Theorem 3.8 (Auslander–Buchsbaum).
Let $(R,\mathfrak{m})$ be a Noetherian local ring and let $M$ be a non-trivial finitely generated $R$-module. If projective dimension of $M$ is finite, then
$\operatorname{projdim}(M)+\operatorname{depth}(M)=\operatorname{depth}(R)$
Proof.
We first establish a preliminary result.W
Proposition 3.9.
For a finitely generated free module $M$ over $R$, for any $R$-module NN, $\operatorname{Ext}_{R}^{i}(N,M)$ vanish if and only if $\operatorname{Ext}_{R}^{i}(N,R)$ vanish.
Proof.
This can be deduced from the fact that $\operatorname{Ext}$ commutes with finite bi-productsW. ∎
As a corollary, $\operatorname{depth}(M)=\operatorname{depth}(R)$ for non-trivial finitely generated free module $M$.W
Returning to the proof of the main theorem, the case $\operatorname{projdim}(M)=0$ is immediate since then $M$ is projective and hence free, the equality follows from the above results.
The case $\operatorname{projdim}(M)=1$ is the core, so we prove it separately, following the proof in [14, Theorem 19.1].
@@
Pick a $k$ basis of $M/\mathfrak{m}M$ and lift it back to $M$, obtaining a map $f:R^{n}\rightarrow M$ whose induced map $(R/\mathfrak{m})^{n}\rightarrow M/\mathfrak{m}M$ is an isomorphism; hence $\mathrm{ker}(f)\subseteq\mathfrak{m}R^{n}$. Since $\mathrm{ker}(f)$ has projective dimension 0, it is projective thus free.
In ordinary proofs, we could view $f$ as a matrix with all elements contained in $\mathfrak{m}$. However, during formalization, viewing a map between free modules as a matrix and operating with it would be complicated, so instead we set up the following two propositions to help.WW
Proposition 3.10.
For $M$ a finitely generated free module over $R$, if the range of $f:M\rightarrow N$ is contained in $IN$ where $I$ is an ideal of $R$, then $f\in Iy\cdot(M\rightarrow N)$
Proposition 3.11.
For $R$-modules $L,M,N$, if $f:M\rightarrow N$ is contained in $\mathrm{Ann}(L)\cdot(M\rightarrow N)$, then the map $\operatorname{Ext}_{R}^{i}(L,M)\rightarrow\operatorname{Ext}_{R}^{i}(L,N)$ induced by $f$ is zero.
Proof.
One verifies this lemma easily by reducing to the case $f=a\cdot g$ with $a\in\mathrm{Ann}(L)$. ∎
Then we know that $i:\mathrm{ker}(f)\hookrightarrow R^{n}$ is contained in $\mathfrak{m}\cdot(\mathrm{ker}(f)\rightarrow R^{n})$, so the second lemma shows that the map $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})$ induced by $i:\mathrm{ker}(f)\hookrightarrow R^{n}$ is zero.
Now consider the long exact sequence induced by $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$:
@@
The first and fourth maps are zero by the discussion above. Moreover, vanishing of $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))$ and of $\operatorname{Ext}_{R}^{i}(k,R^{n})$ are each equivalent to vanishing of $\operatorname{Ext}_{R}^{i}(k,R)$. From the long exact sequence, we deduce that $\operatorname{Ext}_{R}^{i}(k,M)$ vanishes if and only if both $\operatorname{Ext}_{R}^{i}(k,R)$ and $\operatorname{Ext}_{R}^{i+1}(k,R)$ vanish. By the homological characterization of depth, this yields $1+\operatorname{depth}(M)=\operatorname{depth}(R)$.
Finally, we prove the theorem by induction on $\operatorname{projdim}(M)$. Assume the statement holds for projective dimension $n\geq 1$, and let $\operatorname{projdim}(M)=n+1$. Take a surjection $f:R^{m}\rightarrow M$; then $\mathrm{ker}(f)$ has projective dimension $n$. By induction hypothesis, $n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$. Since $\operatorname{depth}(R^{m})=\operatorname{depth}(R)\geq n>0$, so there is no non-trivial $k\rightarrow\mathrm{ker}(f)\hookrightarrow R^{m}$, thus $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$ vanish.
Consider the long exact sequence associated to $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$:
@@
For all $i<\operatorname{depth}(\mathrm{ker}(f))<\operatorname{depth}(R)$, the map $\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))$ is an isomorphism. Together with the vanishing of $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$, we obtain $\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(M)+1$. Hence $n+1+\operatorname{depth}(M)=n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$, which completes the induction and the proof. ∎